OpenStudy (anonymous):

An angles measure is 6 degrees more than 3 times its compliment. find the measure of the angle. a math tutor told me to let the angle be 0 6 degrees + 30 = 90 -6 -6 30/3=84 degrees 0=28 degrees then i check the back of my math book and it said the answer was 69 degrees help?

OpenStudy (anonymous):

or lets make the 0 x so they told me 6 degrees + 3x=90 -6 -6 3x/3=84 degrees x=28 degrees how did the get 28 and my book 69? am i supposed to multiply the 3 instead of divide?

OpenStudy (anonymous):

the answer is 69, the right equation is : 3x+6 = 90-x 4x = 84 x=21 remember that it is 3x plus 6 so it is (3x21)+6 = 69 degrees

OpenStudy (anonymous):

oh ok, that makes sense now. thank you